\begin{thebibliography}{9999}

\bibitem[A]{A} 
 Adler, M., On a trace functional for formal pseudo differential operators and the symplectic structure of the Korteweg-de\thinspace Vries type equations. Invent. Math. 50 (1978/79), no. 3, 219--248.

\bibitem[BL]{BL} 
Bernstein J., Leites D.,  Irreducible representations of type $Q$, odd trace and odd
determinant. C. r. Acad. Bulg. Sci.,
\textbf{35} (1982), no.~3, 285--286

\bibitem[BLLS]{BLLS} 
Bouarroudj S., Lebedev A., Leites
D., Shchepochkina I., Classifications of simple Lie superalgebras in
characteristic $2$. Internat. Math. Res. Not. (2021) 
\url{https://doi.org/10.1093/imrn/rnab265};  
\url{https://arxiv.org/abs/1407.1695}

\iffalse
\bibitem[Ch]{Ch}
Cherednik I., Double affine Hecke algebras. Knizhnik-Zamolodchikov equations,
and Macdonald operators. Internat. Math. Res. Not.  9 (1992), 171--180
\fi

\bibitem[Di]{Di}
Dixmier J.,   \textit{Alg\`ebres enveloppantes}. Paris, Gauthier-Villars (1974) 360 pp.; \textit{Enveloping Algebras}. (Graduate Studies in Mathematics). American Mathematical Society; New edition (1996) 379 pp.


\bibitem[DSB]{DSB}
Duplij S., Siegel W., Bagger J. (eds.) \textit{Concise Encyclopedia
of Supersymmetry and Noncommutative Structures in Mathematics and
Physics}, 2nd edition, Berlin, New York: Springer, 2005, 561pp.



%\bibitem[EG]{EG}
%Etingof P., Ginzburg V.,
%Symplectic reflection algebras, Calogero-Moser space, and deformed Harish-Chandra homomorphism.  Invent. Math. 147 (2002), no. 2, 243--348

\bibitem[Fei]{Fei}
Feigin, B. L. Lie algebras $\mathfrak{gl}(\lambda)$ and cohomology of a~Lie algebra of differential operators. Russian Math. Surveys 43 (1988), no. 2, 169--170



\bibitem[KhM]{KhM}
Khesin, B., Malikov, F. Universal Drinfeld-Sokolov reduction and matrices of complex size. Commun.Math. Phys. 175  (1996) 113--134;  \url{https://arxiv.org/abs/hep-th/9405116}



\bibitem[KS]{KS}
Konstein S. E., Stekolshchik R.,  Klein operator and the numbers of independent traces and supertraces on the superalgebra of observables of rational Calogero model based on the root system. J. Nonlinear Math. Phys.,  20:2  (2013) 295--308 (For a better written version, see
 \url{https://arxiv.org/abs/1212.0508v2})

\bibitem[KT1]{KT1}
Konstein S. E., Tyutin I. V., The number of independent traces and supertraces on symplectic reflection algebras. J. Nonlinear Math. Phys. 21:3  (2014) 308--335; \url{https://arxiv.org/abs/1308.3190}

\bibitem[KT2]{KT2}
Konstein S. E., Tyutin I. V., Traces and supertraces on the symplectic reflection algebras. Theor. Math. Phys. 198 (2019), no. 2, 249--255

\bibitem[KT3]{KT3}
Konstein S. E., Tyutin I. V., Connection between the ideals generated by traces and by supertraces in the superalgebras of observables of Calogero models.  J. Nonlinear Math. Phys., 27:1 (2020), 7--11;  \url{https://arxiv.org/abs/1909.02781}





\bibitem[Lsos]{Lsos}
Leites D. (ed.), \textit{Seminar on supersymmetry v. $1$. Algebra and
Calculus: Main chapters}, (J.~Bernstein, D.~Leites, V.~Molotkov,
V.~Shander), MCCME, Moscow, 2012, 410 pp  \url{https://staff.math.su.se/mleites/books/2011-sos1.pdf} (in Russian; a~version in
English is in preparation but available for perusal)


\bibitem[L]{L}
Leites D.
New simple Lie superalgebras as queerified associative algebras;   
\url{https://arxiv.org/abs/2203.06917}

\bibitem[Ma]{Ma}
Manin, Yu. I., Algebraic aspects of nonlinear differential equations, J. Soviet Math., 11:1 (1979), 1--122

\bibitem[MR]{MR}
Manin, Yu. I.; Radul, A. O. A supersymmetric extension of the Kadomtsev-Petviashvili hierarchy. Comm. Math. Phys. 98 (1985), no. 1, 65--77


\bibitem[Mon]{Mon}
Montgomery S., Constructing simple Lie superalgebras from associative graded algebras. J. Algebra 195 (1997), no. 2, 558--579. 

\bibitem[Mus]{Mus}
 Musson Ian M., Some Lie superalgebras associated to the Weyl algebras. Proc. Amer. Math. Soc. 127 (1999), no. 10, 2821--2827
 
\bibitem[R]{R}
Reiman A. G., Integrable Hamiltonian systems connected with graded Lie algebras, J. Math. Sci., 19:5 (1982), 1507--1545
 
\bibitem[S]{S}
Scott S., \textit{Traces and Determinants of Pseudodifferential Operators}.  Oxford Mathematical Monographs. Oxford University Press, Oxford, (2010) xiv+676 pp.

\bibitem[Sha]{Sha}
Shander V.,  Vector fields and differential equations on supermanifolds, Funct. Anal. Appl., 14:2 (1980), 160--162

\bibitem[Sh]{Sh}
Shander V.,  Complete integrability of ordinary differential equations on supermanifolds. Funct. Anal. Appl., 17:1 (1983), 74--75

\bibitem[Va1]{Va1}
Vasiliev M.A., Extended higher-spin superalgebras and their realizations in terms of quantum operators. Fortsch.Phys. 36 (1988) 33--62

\bibitem[Va2]{Va2}
Vasiliev M.A., Quantization on sphere and high spin superalgebras. JETP Lett. 50 (1989) 374--377

\bibitem[Va3]{Va3}
Vasiliev M.A., Higher spin algebras and quantization on the sphere and hyperboloid. Int. J. Mod. Phys. A 6 (1991) 1115--1135



\end{thebibliography}

\documentclass[12pt]{amsart}

\usepackage{bbm}
\usepackage{hyperref}

\newcommand {\ssecfont} {\normalfont}
\newcommand {\sssecfont} {\normalfont}

\newcommand {\bcdot}   {\mathbin{\hbox{\raise.4ex\hbox{\bf.}}}} % bold \cdot

\newcommand {\ssbegin}[1]
 {\refstepcounter{subsection}
 \def \secno {\gdef \secno {}{\ssecfont
\thesubsection.\hskip 2ex}%
 }%
 \begin{#1}}

\newcommand {\sssbegin}[1]
 {\refstepcounter{subsubsection}
 \def \secno {\gdef \secno {}{\sssecfont
\thesubsubsection.\hskip 2ex}%
 }%
 \begin{#1}}

\newcommand{\ev}{\bar 0}
\newcommand{\od}{\bar 1}
\newcommand{\One}{{\mathbbm{1}}}
\newcommand{\Span}{\text{Span}}
\newcommand{\tr}{\text{tr}}

\def\nopoint#1#2{}

\begin{document}

\title{Supertraces on queerified algebras} %  of observables in the Calogero--Moser model,\\ Vasiliev higher spin algebras,\\ algebras of complex size matrices}

\author{Dimitry Leites${}^{a}$, Irina~Shchepochkina${}^b$}

\address{${}^a$Department of mathematics, Stockholm University, Roslagsv. 101, 
Stockholm, Sweden\\ dimleites@gmail.com\\
${}^b$Independent University of Moscow,
B. Vlasievsky per., d. 11, RU-119 002 Moscow, Russia\\
irina@mccme.ru}

%\begin{history}
%\received{(??, 2009)} \revised{(??, 2009)} \accepted{(??, 2009)}
%\comby{(xxxxxxxxx)}
%\end{history}

%\date{\today}

%--------------------------------------------------------------------
%--------------------------------------------------------------------
%--------------------------------------------------------------------
\begin{abstract} We describe supertraces on ``queerifications'' (see arxiv:2203.06917) of the algebras of matrices of ``complex size'', algebras of observables of Calogero-Moser model, Vasiliev higher spin algebras, and (super)algebras of pseudo-differential operators. In the latter case, the supertraces establish complete integrability of the analogs of Euler equations to be written.
\end{abstract}

\keywords {Simple Lie superalgebra, queerification, trace, supertrace}

\subjclass[2010]{Primary 17B20, 16W55 Secondary 81Q60, 17B70}

\maketitle

%\iffalse
\markboth{\itshape Dimitry Leites\textup{,} Irina Shchepochkina} {{\itshape Traces and supertraces}}
%\fi

\thispagestyle{empty}

The traces on Lie algebras, and even (not odd) supertraces on Lie superalgebras, are  known to be very useful, for example, in representation theory, see, e.g., \cite{DSB} and references therein. The \textit{odd} supertraces are  less known, hence less popular. In Section~\ref{intinv}, we show one of their usages not previously explored: application to the study of integrability of certain hamiltonian dynamical systems.

Inspired by \cite{BLLS}, where Lie algebras are queerified over an algebraically closed ground field of characteristic $p=2$ in order to produce the complete list of simple finite-dimensional Lie superalgebras in characteristic $p=2$, this new method --- queerification --- producing many new simple Lie superalgebras from associative algebras and superalgebras is applied in \cite{L}  over $\mathbb{C}$ to several infinite-dimensional algebras of interest in theoretical physics. A~ number of papers were devoted to description of traces on these algebras, and supertraces on these algebras considered as superalgebras, see \cite{KS, KT1, KT2, KT3}. In this note, we describe the supertraces on the Lie queerifications of these algebras and superalgebras, having added one more type of examples.

\section{Preliminaries}\label{S1}

\subsection{From associative to Lie} Let $\mathbb{K}$ be an algebraically closed ground field of characteristic $p\neq 2$; unless otherwise stated, we consider $\mathbb{K}=\mathbb{C}$. 

Let $A$ be any associative algebra, let $A^L$ be the Lie algebra whose space is $A$ but multiplication is given by the commutator $[a,b]:=ab-ba$ for any $a,b\in A$. 

Let $A$ be a~$\mathbb{Z}/2$-graded algebra; let $p$ denote the parity function. If $A$ a~$\mathbb{Z}/2$-graded  associative algebra, let $A^S$ be the Lie superalgebra whose space is $A$ but the multiplication is given by the supercommutator, which by the modern habitual abuse of notation is also denoted $[a,b]$ although defined  to be $[a,b]:=ab-(-1)^{p(a)p(b)}ba$. Let $\mathfrak{g}':=[\mathfrak{g},\mathfrak{g}]$ be the first derived Lie algebra (resp. Lie superalgebra), a.k.a. commutant  (resp. supercommutant), of the Lie algebra (resp. Lie superalgebra) $\mathfrak{g}$. 

Not every $\mathbb{Z}/2$-graded algebra $\mathcal{A}$ is called \textit{superalgebra}: only if  multiplication in $\mathcal{A}$ or (if $\mathcal{A}$ is associative) in $\mathcal{A}^S$ depends on the parity. 

Recall, that  a \textit{trace} (called \textit{supertrace}  in the super setting, for emphasis) on a given Liealgebra (resp. Lie superalgebra) $\mathfrak{g}$ is  a linear function that vanishes on its commutant (resp. supercommutant), so there are $\dim (\mathfrak{g}/[\mathfrak{g},\mathfrak{g}])$  linearly independent traces (resp. supertraces) on~ $\mathfrak{g}$, some of the supertraces can even and some of them odd.

\subsection{Queerifications in characteristic  $p\neq 2$ (from \cite{BLLS})}\label{111} Let $A$ be an
associative algebra.  The space of the associative algebra $\text{Q}(A)$ --- the \textit{\textbf{associative} queerification} of $A$ --- is $A \oplus\Pi(A)$, where
$\Pi$ is the change of parity functor, with the same multiplication in $A$ and the adjoint action of  $A$ on $\Pi(A)$; let 
\[
\Pi(x)\Pi(y):= xy\text{~for any~~}x,y\in A.
\]
Set $\text{Q}(n):=\text{Q}(\text{Mat}(n))$. 

We will be mostly interested in the following \textit{\textit{\textbf{Lie}} queerifications} $\mathfrak{q}(A)$:  

1) for $A$ an associative algebra;

2) for $A$ a $\mathbb{Z}/2$-graded associative algebra (\lq\lq Liefication" yields $\mathbb{Z}/2$-graded Lie algebra $A^L$);

3) for $A$ an associative superalgebra (this case differs from case 2) because passing to  $A^S$ the supercommutators instead of commutators are considered; \lq\lq super Liefication" yields Lie superalgebra $A^S$ which is $\mathbb{Z}/2$-graded  by parity. %; considered as algebra it is $(\mathbb{Z}/2\times\mathbb{Z}/2)$-graded).

In case 1), as spaces, $\mathfrak{q}(A):={A^L\oplus\Pi(A)}$, so $\mathfrak{q}(A)_{\bar 0}=A^L$ and $\mathfrak{q}(A)_{\bar 1}=\Pi(A)$, with the bracket given
by the following expressions and super anti-symmetry, i.e., anti-symmetry amended by the Sign Rule:
\begin{equation}\label{commRel}
{}[x,y]:=xy-yx;\quad [x,\Pi(y)]:=\Pi(xy-yx);\quad [\Pi(x),\Pi(y)]:=
xy+yx \ \text{~for any~~}x,y\in A.
\end{equation}

The term ``queer", now conventional, is taken after the Lie
superalgebra $\mathfrak{q}(n):=\mathfrak{q}(\text{Mat}(n))$. (The associative superalgebra $\text{Q}(n)$ is an analog of $\text{Mat}(n)$; likewise, the Lie superalgebra $\mathfrak{q}(n)$ is an analog of $\mathfrak{gl}(n)$  for several reasons, for example, due the role of these analogs in Schur's Lemma and in the classification of central simple superalgebras, see \cite[Ch.7]{Lsos}.) We express the elements
of the Lie superalgebra $\mathfrak{q}(n)$ by means of a~pair of matrices
\begin{equation}
\label{.1} (X,Y)\longleftrightarrow
\begin{pmatrix}X&Y\\Y&X
\end{pmatrix}
\in \mathfrak{gl}(n|n), \text{~~where $X,Y \in \text{Mat}(n)$}.
\end{equation}

For any associative $A$, we will similarly denote the elements of $\mathfrak{q}(A)$ by pairs $(X,Y)$, where $X,Y \in A$. The brackets between these elements are as follows:
\begin{equation}\label{.2}
\renewcommand{\arraystretch}{1.4}
\begin{array}{l}
{}[(X_1,0),(X_2,0)]:=([X_1,X_2],0),\quad
[(X,0),(0,Y)]:=(0,[X,Y]),\\
{}[(0,Y_1), (0,Y_2)]:=(Y_1Y_2+Y_2Y_1,0).
\end{array}\end{equation}

We define Lie queerifications in cases 2) and 3) in the next section.

\section{Traces and supertraces: generalities}{}~{}

\ssbegin{Theorem} Let $A$ be an associative algebra. Then, the estimate of the number of traces in the $3$ cases  of its Lie queerification are described below. \end{Theorem}

\subsubsection{Case $1$}\label{2.1}
Let $A$ be an associative algebra, then $Q(A):=A\oplus \Pi(A)$ is the \textit{associative queerification} of $A$; let $\mathfrak{g}:=A^L$ be the corresponding Lie algebra and let the Lie superalgebra $\mathfrak{q}\mathfrak{g}:=(Q(A))^S=\mathfrak{q}(A)$ be the \textit{Lie queerification} of $A$.
Clearly, $\mathfrak{q}\mathfrak{g}=\mathfrak{g}\oplus \Pi(\mathfrak{g})$ as spaces.

Denote, for brevity, $\mathfrak{u}:=(\mathfrak{q}\mathfrak{g})'$. By definition the supercommutant $\mathfrak{u}:=\mathfrak{u}_{\ev}\oplus \mathfrak{u}_{\od}$ is spanned by elements $[a,b]$ for any $a,b\in \mathfrak{q}\mathfrak{g}$. In particular, $\mathfrak{u}_{\od}$ is spanned by elements of the form $[x,\Pi(y)]=\Pi([x,y])$ for any $x,y\in  (\mathfrak{q}\mathfrak{g})_{\ev}$, and hence $\mathfrak{u}_{\od}=\Pi([\mathfrak{g},\mathfrak{g}])$. Therefore, there are as many odd supertraces on $\mathfrak{q}\mathfrak{g}$ as there are traces on $\mathfrak{g}$.

Clearly, $\mathfrak{u}_{\ev}$ is the sum of two ideals of $\mathfrak{g}$:
\[
\mathfrak{u}_{\ev}=[\mathfrak{g},\mathfrak{g}]+[\Pi(\mathfrak{g}),\Pi(\mathfrak{g})].
\]
Observe that the second summand does not have to be contained in the first one. Therefore, there are fewer even supertraces on $\mathfrak{q}\mathfrak{g}$ than there are traces on $\mathfrak{g}$.

In particular, if  $A$ has unit $\One$, 
%and $\mathfrak{g}=\mathbb{K}\cdot \One\oplus \mathfrak{g}'$,  
then $\mathfrak{u}_{\ev}=\mathfrak{g}$, because $[\Pi(\One),\Pi(x)]=2x$ for any $x\in \mathfrak{g}_{\ev}$.

For example, if $A=\text{Mat}(n)$, then on $Q(A)$ there is an odd trace, nowadays called \textit{queertrace}; it was first defined in \cite{BL} by the formula
\[
\text{qtr}: \begin{pmatrix}X&Y\\Y&X
\end{pmatrix}\mapsto \tr Y.
\]

\subsubsection{Case $2$}\label{2.3}
Let $A$ be a $\mathbb{Z}/2$-graded associative algebra, $\mathfrak{g}:=A^L$. Clearly, $\mathbb{Z}/2$-grading of~ $A$ makes $\mathfrak{g}=\mathfrak{g}_{\ev}\oplus \mathfrak{g}_{\od}$ a~ $\mathbb{Z}/2$-graded Lie algebra.

Now, consider $A$ as an associative superalegbra; let $\mathfrak{s}=\mathfrak{s}_{\ev}\oplus \mathfrak{s}_{\od}:=A^S$. Clearly, $\mathfrak{s}_{\ev}=\mathfrak{g}_{\ev}$ and $\mathfrak{s}_{\od}=\Pi(\mathfrak{g}_{\od})$, as spaces. The first equality is, moreover, true as an isomorphism of Lie algebras, whereas the second one as an isomorphism of  $\mathfrak{s}_{\ev}=\mathfrak{g}_{\ev}$-modules. 

The supercommutants are as follows:
\[
\begin{tabular}{|l|}
\hline
$\mathfrak{g}'=(\mathfrak{g}')_{\ev}\oplus (\mathfrak{g}')_{\od}$,  where for any $x,y\in\mathfrak{g}_{\ev}$ and $a,b\in\mathfrak{g}_{\od}$, we have\\ 
$(\mathfrak{g}')_{\ev}=\Span([x,y]=xy-yx,[a,b]=ab-ba),\ (\mathfrak{g}')_{\od}=\Span([x,a]=xa-ax)$;\\
\hline
$\mathfrak{s}'=(\mathfrak{s}')_{\ev}\oplus (\mathfrak{s}')_{\od}$,  where for any  $x,y\in\mathfrak{s}_{\ev}$, $a,b\in\mathfrak{s}_{\od}$ we have\\
$(\mathfrak{s}')_{\ev}=\Span([x,y]=xy-yx,[a,b]=ab+ba), \ (\mathfrak{s}')_{\od}=\Span([x,a]=xa-ax)$.\\
\hline
\end{tabular}
\]

Therefore, $\mathfrak{s}'_{\od}=\Pi((\mathfrak{g}')_{\od})$, but \text{a priori}  we can not say anything about relations between $(\mathfrak{g}')_{\ev}$ and $(\mathfrak{s}')_{\ev}$, except for both contain $(\mathfrak{g}_{\ev})'=(\mathfrak{s}_{\ev})'$.

\subsubsection{Case $3$}\label{2.2}
Now, let $A=A_{\ev}\oplus A_{\od}$ be an associative superalgebra, $\mathfrak{g}=\mathfrak{g}_{\ev}\oplus \mathfrak{g}_{\od}:=A^S$. 

Let $Q(A):=A\oplus \Pi(A)$ and $\mathfrak{q}\mathfrak{g}:=(Q(A))^S$; clearly, $\mathfrak{q}\mathfrak{g}=\mathfrak{g}\oplus \Pi(\mathfrak{g})$ as superspaces.
It is also clear that the natural $(\mathbb{Z}/2\times \mathbb{Z}/2)$-grading on $Q(A)$  induces the same grading on  $\mathfrak{q}\mathfrak{g}$, where
$$
(\mathfrak{q}\mathfrak{g})_{\ev,\ev}=\mathfrak{g}_{\ev}, \; (\mathfrak{q}\mathfrak{g})_{\ev,\od}=\mathfrak{g}_{\od}, \; (\mathfrak{q}\mathfrak{g})_{\od,\ev}=\Pi(\mathfrak{g}_{\od}), \; (\mathfrak{q}\mathfrak{g})_{\od,\od}=\Pi(\mathfrak{g}_{\ev}).
$$

Let us now compare the supercommutants of $\mathfrak{g}$ and $\mathfrak{q}\mathfrak{g}$. We will denote the elements of $\mathfrak{g}_{\ev}$ by letters $x, y,\dots$, the elements of $\mathfrak{g}_{\od}$ by letters $a,b,c,\dots$.
We have $\mathfrak{g}'=(\mathfrak{g}')_{\ev}\oplus (\mathfrak{g}')_{\od}$,  where 
$$
 (\mathfrak{g}')_{\ev}=\Span([x,y]:=xy-yx, \ \ [a,b]:=ab+ba), \ \ (\mathfrak{g}')_{\od}=\Span([x,a]:=xa-ax).
$$

The elements that span homogeneous components of $\mathfrak{q}\mathfrak{g}$ are as follows:
\[
\begin{tabular}{|c|c|c|c|}
\hline
$\mathfrak{u}_{\ev,\ev}$ & $\mathfrak{u}_{\ev,\od}$ & $\mathfrak{u}_{\od,\ev}$ & $\mathfrak{u}_{\od,\od}$\\
\hline
$[x,y]=xy-yx$ & $[x,a]$ & $[x,\Pi(a)]$ & $[x,\Pi(y)]$\\
$[a,b]=ab+ba$ &$=xa-ax$ &$=\Pi(xa-ax)$ &$=\Pi(xy-yx)$\\
$[\Pi(a),\Pi(b)]=ab-ba$ & $[\Pi(a),\Pi(x)]$ & $[a,\Pi(x)]$ & $[a,\Pi(b)]$\\
$[\Pi(x),\Pi(y)]=xy+yx$ &$=ax-xa$ &$=\Pi(ax+xa)$ &$=\Pi(ab-ba)$ \\
\hline
\end{tabular}
\]

Therefore, $\mathfrak{u}_{\ev,\ev}=(A_{\ev})^2+(A_{\od})^2$ and $\mathfrak{u}_{\od,\ev}=\Pi(A_{\ev} \cdot A_{\od})$, as spaces. In particular, if $A$ has unit,  then 
\[
\text{$\mathfrak{u}_{\ev,\ev}=(\mathfrak{q}\mathfrak{g})_{\ev,\ev}=\mathfrak{g}_{\ev}$ and $\mathfrak{u}_{\od,\ev}=(\mathfrak{q}\mathfrak{g})_{\od,\ev}=\Pi(\mathfrak{g}_{\od})$.}
\]

In general, \text{a priori} we can only say that 
$\mathfrak{u}_{\ev,\od}=(\mathfrak{g}')_{\od}$.

\section{Examples of supertraces on queerified algebras and superalgebras}

\subsection{Clifford--Weyl algebras and superalgebras}\label{3.1} Vasiliev was, most probably, the first to publish that \textit{on the Weyl  algebra $W_n$ of polynomial differential operators in $n$ even indeterminates $x_i$ considered as superalgebra with parity given by $p(x_i)=p(\partial_i)=\od$ for all $i$, there is an even supertrace}, see \cite{Va1}. For a generalization to the Calogero-Moser case, see \cite{Va2}; the detailed version \cite{Va3} contains an elementary proof of uniqueness of the (super)trace; the construction yields algebras of \lq\lq matrices of complex size" which first appeared as associative algebras in the book \cite{Di} and as Lie algebras in \cite{Fei}.
For a~ generalization to symplectic reflection algebras  in \cite[Th.7.1.1]{KT1}. Alexey Lebedev suggested a~ beautiful elementary proof of the existence of the trace on $W_n$, see \S\ref{sL}. 

Obviously unaware of Vasiliev's works, his results were rediscovered by mathematicians, see \cite[Proposition~4.3]{Mon} and \cite{Mus}. Montgomery found out the following sufficient condition to the super version of Herstein's theorem and formulated it in the infinite-dimensional situation (in the finite-dimensional case, it is also true). 

\sssbegin{Theorem}{\textup{(\cite[Th.3.8]{Mon})}}\label{ThM} Let $A$ be a~$\mathbb{Z}/2$-graded simple associative algebra of characteristic $p\neq 2$ with supercenter $Z$ whose elements supercommute with any $a\in A$. Let the condition
\begin{equation}\label{Cond}
\text{if $u^2 \in Z$, then $u\in Z$ for any homogeneous $u\in A_{\bar 1}$}
\end{equation} 
hold. Then, $SL(A):=(A^S)^{(1)}/((A^S)^{(1)}\cap Z)$ 
is a~simple Lie super algebra.\end{Theorem}

Observe that the superalgebra $A$ of differential operators in any finite number of  indigenously odd indeterminates (a.k.a. the \textit{Clifford} algebra on $2n$ generators considered as a $\mathbb{Z}/2$-graded associative superalgebra) is isomorphic to the matrix superalgebra $\text{Mat}(2^{n-1}|2^{n-1})$ on which there is an even supertrace, whereas on $\mathfrak{q}(2^{n}):=\mathfrak{q}(\mathfrak{gl}(2^{n-1}|2^{n-1}))$ there is the (well-known today) odd \textit{queer trace}.
Therefore, by arguments in Subsection~\ref{2.3}, \textbf{there are no even supertraces on $\mathfrak{q}A$, but there is one odd supertrace}.

\subsection{(Super)algebras of ``matrices of complex size''}\label{3.2} Absolutely the same arguments as in Example \ref{3.1} are applicable to the queerifications of algebras and superalgebras $A$ of ``matrices of complex size'', see \cite[Subsection 2.2]{L} with the same  answer. (Although we do not need it here, note --- just for its beauty --- that since $\mathfrak{gl}(\lambda)=\mathfrak{gl}(\lambda)'\oplus \mathbb{C}\,1$, one can define the trace on the Lie algebra $\mathfrak{gl}(\lambda)$ by any value on~ $1$. J.~Bernstein defined the trace on $\mathfrak{gl}(\lambda)$ for $\lambda\in\mathbb{C}\setminus\mathbb{Z}$, see \cite{KhM}, so that $\tr(1)=\lambda$; Bernstein's trace naturally generalizes the trace on $\text{Mat}(|n|)$  for any $n\in\mathbb{Z}\setminus\{0\}$ such that $\tr(1_{|n|})=|n|$.)

\subsection{Symplectic reflection algebras and superalgebras}\label{3.3} For the classification of traces (resp. supertraces) on these algebras and superalgebras $A$, see \cite[Tables on pp.5,6]{KS}. Considering them as algebras (resp. superalgebras) we get the exact number of supertraces on their Lie qeerifications, according to general results established in Case \ref{2.1} (resp. Case \ref{2.3}).

For a description of ideals in these algebras and superalgebras, see \cite{KT1,KT2, KT3}. It is interesting to determine when these ideals are themselves simple algebras and superalgebras, and describe (super)traces on them and on their queerifications.

\subsection{(Super)trace on the (super)algebra of pseudo-differential operators}\label{SSpseu}
 
 \subsubsection{$N=0$} Recall that the associative algebra $\Psi$ of pseudo-differential operators of integer order is $\mathcal{F}((D^{-1}))$, where $D:=\frac{d}{dx}$  and $\mathcal{F}$ is the algebra of functions in $x$, with multiplication given for any integer $n$ by the Leibniz rule 
 \[
 D^{n}f:=\sum_{k\geq 0}\binom{n}{k}D^{k}(f)D^{n-k}, \text{~~where  $\binom{n}{k}:=\frac{n(n-1)\dots(n-k+1)}{k!}$, for any $f\in\mathcal{F}$}.
 \]
Adler defined a~ trace on the algebra $\Psi$ of pseudo-differential operators, see \cite{A} and a very reader-friendly reviews \cite{R, Ma}, as the composition of the residue and the indefinite integral
\[
\tr(\sum_{k\leq n}f_kD^{k})=\int f_{-1}dx, \text{~~where $f_k\in\mathcal{F}$}. 
\]
This trace (it vanishes on the commutators even before the integral is taken, just residue suffices, see \cite{R}) takes values in $\mathcal{F}$. According to Subsection~\ref{2.3}, \textbf{there are no even supertraces on $\mathfrak{q}\Psi$, but there are $\geq 1$ odd supertraces}; we conjecture there is just one odd supertrace.
 
\subsubsection{$N=1$}\label{ss3.4.2} On the superalgebra $\Psi_{1}:=\mathcal{F}((D^{-1}))$ of $N=1$-extended pseudo-differential operators, where $D:=\frac{\partial}{\partial \xi}+\xi\frac{\partial}{\partial x}$ and $\mathcal{F}$ is the algebra of functions in the even $x$ and odd~ $\xi$, Manin and Radul defined an even supertrace, see \cite{MR}. According to Subsection~\ref{2.3}, \textbf{there are no even supertraces on $\mathfrak{q}\Psi_{1}$, but there are $\geq 1$ odd supertraces}; we conjecture there is just one odd supertrace.

\section{An application of traces: integrals in involution}\label{intinv} Let $A$ be an associative (super)algebra, $\tr$ a (super)trace on~ $A$, and  $b$ the corresponding  \textit{invariant symmetric} bilinear form (briefly: IS form)
\[
b(X, Y):=\tr(XY)\text{~~for any $X, Y\in A$}. 
\]
Let, moreover, $b$ be non-degenerate, briefly: NIS. 

Let $\tr$, and hence $b$, be even. Then, for the equation ($L$ and $P$ are in honor of Peter Lax)
\begin{equation}\label{Lax}
\dot L=[L, P], \text{~~where $L, P\in A$ and dot signifies derivative with respect to time $t$},
\end{equation}
and for its more subtle version, more adequate in the super setting with $1|1$-dimensional Time with even coordinate $t$ and odd one $\tau$, see \cite{Sha}:
\begin{equation}\label{11Lax}
(\partial_\tau+\tau\partial_t) L=[L, H], \text{~~where $L, H\in A$},
\end{equation}
the functions $L\mapsto \tr(L^k)$ are integrals in involution with respect to the Poisson bracket defined on the space of functions on $\mathfrak{g}^*$ as follows, see, e.g., \cite{A}. We identify $\mathfrak{g}$ with the linear functions on $\mathfrak{g}^*$ and for any functions $f,g\in\mathcal{F}(\mathfrak{g}^*)$ the Poisson bracket is defined to be
\begin{equation}\label{PoiB}
\{f,g\}(X):=X([df(X),dg(X)])\text{~~for any $X\in\mathfrak{g}^*$}.
\end{equation}

It seems, nobody considered yet the Euler equations or Lax pairs \eqref{Lax} related with superalgebras $A$ on which there is an odd trace $\text{qtr}$, and hence an odd $b$. If $b$ is odd, then $\mathfrak{g}\simeq\Pi\mathfrak{g}^*$, and an antibracket, rather than a~Poisson bracket, is defined on the space $\Pi\mathcal{F}(\mathfrak{g}^*)$ of functions on $\mathfrak{g}^*$. The functions $L\mapsto\text{qtr}(L^k)$ are integrals in involution, i.e., commuting with the Hamiltonian and each other (themselves including in the case of the anti-bracket) with respect to the antibracket.

On $2n|k$-dimensional superspace on which the Poisson bracket is defined or $n|n$-dimensional superspace  which the anti-bracket bracket is defined, let there be $n$ first integrals of the dynamical system~ \eqref{Lax} in involution. Then, a theorem of Shander guarantees complete integrability of the system whatever $k<\infty$ is, see \cite{Sh}. 

The traces on (super)algebras $A$ considered in Subsection~\ref{SSpseu} determine what researchers conceded to call ``complete integrability" in the case of infinite-dimensional Hamiltonian system (since there are infinitely many of these traces --- \lq\lq this infinity is a half of the infinite dimension"). Examples: the KdV equations (resp. $(N=1)$-extended KdV, see Subsection~\ref{ss3.4.2}) for the case where $L$ is the Schr\"odinger operator (resp. its $(N=1)$-extension). 

\textbf{Open problem}: For any simple associative algebra $A$, give examples of integrable systems related with $Q(A)$.

\section{Supertrace on $W_n$}\label{sL}

\subsection{Two general facts (\cite{S})}\label{GF}

1) If there is no (super)trace on an associative (super)algebra $A$, then there is no (super)trace on any product $A\otimes B$ for any associative (super)algebra $B$ with unit: if $[a_1, a_2] = a$, then  $[a_1\otimes 1, a_2\otimes b] = a\otimes b$ for any $a_1, a_2\in A$ and $ b\in B$, i.e., if any element of  $A$ can be represented as a linear combination of (super)commutators, then the same is true for any element of  $A\otimes B$.
 
2) Let $\tr_i$ be a  (super)trace on the associative (super)algebra $A_i$ for $i=1,2$. Then,  
\[
\text{$\tr (a_1\otimes a_2) := (\tr_1\, a_1)(\tr_2\,  a_2)$ for any $a_i\in A_i$}
\]
 is a (super)trace on $A_1\otimes A_2$.

\ssbegin{Lemma}\label{Le} Consider $W_n$ as a superalgebra with $p(x_i)=p(\partial_{x_i})=\od$. 

Then, on $W_n$, there is an even trace.\end{Lemma}

Observe that if we consider $W_n$ as an algebra, not a~superalgebra, no analog of Lemma~\ref{Le} takes place since the associative algebra $W_n$ is simple; on the other hand, the center (constants) is given by a non-trivial cocycle.

\begin{proof}(A. Lebedev) Actually, $W_n=\mathbb{K}1\oplus [W_n, W_n]$, where $[W_n, W_n]$ is the supercommutant. The proof below demonstrates existence of the trace, but its uniqueness (up to a~non-zero factor) should be proved separately. For the proof of uniqueness, see \cite{Va3, Mon}.

Let $n=1$. Let $d:=\frac{d}{dx}$, let the \text{weight} $\text{wht}$ of $x$ be equal to $1$, let $\text{wht}(d):=-1$. On $W_n$, define the following linear function $T$:
\[
T(P) :=\begin{cases} (P(\frac{1}{x+1}))|_{x=1}&\text{ if $\text{wht}(P)=0$,}\\
0&\text{ if $\text{wht}(P)\neq 0$}.
\end{cases}
\]
Let us prove that $T$ is a supertrace on $W_n$, i.e.,
\[
T(PQ) = (-1)^{p(P)p(Q)}T(QP). 
\]
Clearly, it suffices to prove this for the case where $P$ and $Q$ are monomials whose weights are opposite.
 
Case 1: $P= x^{n+1}d^n$ and $Q=d$. Then,
\[
\begin{array}{ll}
T(PQ) &= (x^{n+1} (-1)^{n+1} \frac{(n+1)!}{(x+1)^{n+2}})|_{x=1} = -(-\frac12)^{n+2} (n+1)! \ ,\\ 
T(QP) &= (d (x^{n+1} (-1)^n \frac{n! }{(x+1)^{n+1}})|_{x=1} \\
&= ( (-1)^n \frac{(n+1)! \ x^n }{(x+1)^{n+1}} - (-1)^n \frac{(n+1)! \ x^{n+1}}{(x+1)^{n+2}})|_{x=1} = (-\frac12)^{n+2} (n+1)!
\end{array}
\]
 
This implies the answer for the case where $Q=d$ and any $P$ of weight $1$, because any such $P$ can be represented as a linear combination of operators of the form $x^{n+1}d^n$.
 
Case 2: $\text{wht}(P)=-1$ and $Q=x$. Then,
\[
\begin{array}{l}
T(QP) = (xP(\frac{1}{x+1}))|_{x=1} = (P(\frac{1}{x+1}))|_{x=1},\\
T(PQ) = (P(\frac{x}{x+1}))|_{x=1} = (P(1-\frac{1}{x+1}))|_{x=1} = -(P(\frac{1}{x+1}))|_{x=1}, 
\end{array}\]
since $P(1)=0$ because $\text{wht}(P)<0$.
 
This implies the general case where  $P$ and $Q$ are monomials of opposite weights, because we can transplant $x$ and $d$, one by one, from the end of $PQ$ to the beginning until we get $QP$, and each transplantation changes the sign by the opposite;  by $(-1)^{deg(Q)} = (-1)^{p(P)p(Q)}$ altogether.
 
Since $T(1)=\frac12$, it follows that 1 can not be represented as a linear combination of supercommutators. 

For $n>1$, recall that $W_n\simeq W_{n-1}\otimes W_1$ and apply general fact 2), see Subsection~\ref{GF}. \end{proof} 

\textbf{Comment: how to guess the form of $T$}. Let $P$ be a differential operator of weight~ 0.  In the basis $1$, $x$, $x^2$, \dots , consider $P$ as a linear operator on the space of polynomials and consider the matrix of $P$. Naively, ignoring possible divergence, the supertrace of this matrix is equal to
\[
\sum_{n=0}^\infty (-1)^n \text{(the coefficient of $x^n$ in $Px^n$)}.
\]
Since $\text{wht}(P)=0$, then $Px^n$ is equal to the above-mentioned coefficient of $x^n$. In other words, the coefficient is equal to the value of $Px^n$ at $x=1$. Hence, the supertrace is equal to
\[
\sum_n (-1)^n Px^n|_{x=1} = P(1-x+x^2-x^3+...)|_{x=1} = \left(P\left(\frac{1}{1+x}\right)\right)|_{x=1}.
\]

\subsection*{Acknowledgements} We are thankful to A.~Lebedev for help. DL was supported by the grant AD 065 NYUAD.

%--------------------------------------------------------------------
%--------------------------------------------------------------------
%--------------------------------------------------------------------
\end{document}